1/3y + 3 = 8- 1/6y
Answer by edjones(8007)

(1/3)y + 3 = 8- (1/6)y
2y+18=48-y Multiply each side of the equation by 6.
3y=30 Add y-18 to each side.
y=10
.
Ed